OpenStudy (anonymous):

6 (f+1) -3 (f+2) = 3 (2f+3) +6

OpenStudy (whpalmer4):

Use the distributive property to expand everything: \(a(b+c) = ab + ac\) Then add and subtract terms from each side to collect all of the f's on one side, and a number on the other.

OpenStudy (anonymous):

im still confused can u help me solve it

OpenStudy (whpalmer4):

Okay, 6(f+1) = 6f + 6(1). Do that for the rest of the equation. Be careful: -3(f+2) = -3f + (-3)(2) = -3f -6.

OpenStudy (anonymous):

so you are going to use foil. you will get 6f+6-3f-6=6f+9+6 then you combine to solve for "f"

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

6f + 6 -3f - 6 = 6f + 9 +6

OpenStudy (whpalmer4):

Okay, now combine like terms. For example, the left side would be 6f - 3f + 6 - 6 which simplifies to 3f.

OpenStudy (anonymous):

6(f+1)-3(f+2)=3(2f+3)+6 Multiply 3 by each term inside the parentheses. 6(f+1)-3(f+2)=6f+9+6 Add 6 to 9 to get 15. 6(f+1)-3(f+2)=6f+15 Multiply 6 by each term inside the parentheses. 6f+6-3(f+2)=6f+15 Multiply -3 by each term inside the parentheses. 6f+6-3f-6=6f+15 Since 6f and -3f are like terms, add -3f to 6f to get 3f. 3f+6-6=6f+15 Subtract 6 from 6 to get 0. 3f=6f+15 Since 6f contains the variable to solve for, move it to the left-hand side of the equation by subtracting 6f from both sides. 3f-6f=15 Since 3f and -6f are like terms, add -6f to 3f to get -3f. -3f=15 Divide each term in the equation by -3. -(3f)/(-3)=(15)/(-3) Simplify the left-hand side of the equation by canceling the common factors. f=(15)/(-3) Simplify the right-hand side of the equation by simplifying each term. f=-5